# Compositional Zero-shot Learning via Progressive Language-based Observations

### v1

@@ Abstract
Compositional zero-shot learning aims to recognize unseen state-object compositions by leveraging known primitives (state and object) during training. However, effectively modeling interactions between primitives and generalizing knowledge to novel compositions remains a perennial challenge. There are two key factors: object-conditioned and state-conditioned variance, i.e., the appearance of states (or objects) can vary significantly when combined with different objects (or states). For instance, the state “old” can signify a vintage design for a “car” or an advanced age for a “cat”. In this paper, we argue that these variances can be mitigated by predicting composition categories based on pre-observed primitive. To this end, we propose Progressive Language-based Observations (PLO), which can dynamically determine a better observation order of primitives. These observations comprise a series of concepts or languages that allow the model to understand image content in a step-by-step manner. Specifically, PLO adopts pre-trained vision-language models (VLMs) to empower the model with observation capabilities. We further devise two variants: 1) PLO-VLM: a two-step method, where a pre-observing classifier dynamically determines the observation order of two primitives. 2) PLO-LLM: a multi-step scheme, which utilizes large language models (LLMs) to craft composition-specific prompts for step-by-step observing. Extensive ablations on three challenging datasets demonstrate the superiority of PLO compared with state-of-the-art methods, affirming its abilities in compositional recognition.
1 Introduction
“There’s more than one way to skin a cat.”
@@ 1 Introduction
What enables us humans to recognize new concepts we have never encountered before? It all comes down to our capacity to generalize learned knowledge to unseen domains. For instance, when presented with concepts “green apple” and “yellow banana”, we can instantly recognize and imagine the concept “green banana” by combining state “green” with object “banana”. Inspired by this innate cognitive ability of humans, Compositional Zero-Shot Learning (CZSL) emerges to tackle the challenge of recognizing unseen state-object compositions (e.g., “green banana”) by leveraging visible primitives (i.e., state and object) in compositional concepts during training and applying the knowledge during inference [25, 29, 19].
@@
Figure 1: Illustrations of challenges in CZSL and the proposed PLO method. (a) The challenge of object/state-conditioned variation: A perceptible variance emerges in the visual appearance of state/object primitives when juxtaposed in different compositions. (b) PLO-VLM/LLM: A two/multi-step observation approach dynamically controls the observation order for effective recognition.
@@
Effectively modeling the interactions between state and object primitives, as well as extrapolating the understanding of seen compositions to unseen ones, poses major challenges in CZSL. Concretely, it revolves around two critical factors: 1) Object-conditioned Variance: Wherein the visual representations of the same state category can vary considerably when different objects are involved. As depicted in Figure 1(a), considering the state “old” in the context of modifying a “car” and a “cat”, it may refer to a vintage design with classic curves and retro elements for the “car”, evoking a sense of nostalgia and history, whereas for the “cat”, it denotes the senior age of feline with features like grey fur, reflecting the passage of time and aging process. 2) State-conditioned Variance: It pertains to the variations in the appearance of an object when combined with different states. As shown in Figure 1(a), for composition “peeled banana”, the “banana” exhibits a smooth texture and a pale appearance, as the outer peel is removed. In contrast, for the composition “sliced banana”, the “banana” takes on a sliced appearance with visible segments.
Previous approaches in CZSL often construct separate classifiers for recognizing states and objects simultaneously, overlooking their intrinsic relationship. Recent efforts have made strides in addressing the first factor by adopting a two-stage method with an object-then-state order [36, 11, 15]. Prioritizing the prediction of the object primitive allows the model to capture salient visual cues (e.g., shapes), thereby enhancing the overall comprehension of compositions. Subsequently, armed with the knowledge of the object primitives, the CZSL model sequentially refines its understanding by classifying the state primitives conditioned on guided object features.
@@
Nonetheless, we argue that there is more than one way to skin a cat, and the human cognition process will progressively collect different observations for specific compositions, in a simple to complex manner. In certain cases, such as the composition “ripe banana” in Figure 1(b), the object itself, “banana”, possesses highly salient visual cues that make it easily recognizable due to its curving shape and vibrant yellow color. Once we establish that it is a “banana”, we can then further analyze its state and recognize it as a “ripe banana” by observing additional visual cues, e.g., the presence of brown spots on the yellow skin. In contrast, for compositions like “mashed banana”, possess distinct visual features primarily related to the state “mashed” rather than the object. The mushy texture becomes the prominent aspect that captures our attention. Consequently, through further analysis of extra visual features, e.g., yellow and sticky material, we refine our recognition and discern it as a “mashed banana”.
@@
In this paper, inspired by the human-like cognition process, we propose a novel approach, Progressive Language-based Observations (PLO) for CZSL. Specifically, PLO dynamically determines the order of progressive observations in the form of language, building upon the pre-trained vision-language models (VLMs), e.g., CLIP [34]. These observations comprise a series of languages that allow the model to observe the image’s content step by step. Due to training with image-text pairs, these VLMs endow the model with observing capability by measuring the similarity between the two modalities within the same space. For dynamic progressive observation, we propose two variants: PLO-VLM and PLO-LLM. In PLO-VLM, we introduce a two-step observation strategy that adopts a pre-observing classifier based on VLM to dynamically determine the order of primitive-containing languages based on image features. Subsequently, leveraging the observed primitive knowledge (semantic features from primitive prompts), we integrate this information via a cross-modal attention module for category prediction of the remaining primitive. In PLO-LLM, we further extend to a multi-step observation scheme, employing large language models (LLMs), e.g., GPT [4] to design composition-specific prompts (e.g., “yellow, mushy substance” in Figure 1(b)) for each composition category, thus obtaining a composition-specific observation order. This method allows us to selectively extract features at each observation step, boosting the model’s ability to understand and recognize composition categories more effectively.
@@
Three popular and challenging CZSL datasets MIT-States [12], UT-Zappos [29], and C-GQA [39] are used for evaluation. Extensive results show that our PLO exceeds the current state-of-the-art CZSL methods with significant gains in both closed-world and open-world settings. In summary, the main contributions of our work are three-fold:
  • We propose the novel PLO for CZSL. To the best of our knowledge, it is the first work to dynamically allocate the order of observations using language, enabling effective prediction of unseen state-object compositions.
@@
  • We introduce two model variants: 1) PLO-VLM: It uses a pre-observing classifier to determine observation order based on image features. 2) PLO-LLM: It employs LLMs to design composition-specific observation prompts.
@@
  • Extensive results on multiple datasets demonstrate that PLO outperforms existing methods, showcasing its effectiveness in recognizing unseen state-object compositions.
2 Related Work
@@
Compositional Zero-Shot Learning (CZSL). CZSL is a prominent research area, and state-of-the-art methods can be broadly categorized into two directions: 1) Composed CZSL [28, 33, 29, 2, 26]: It focuses on directly classifying compositional state-object, which projects whole compositional visual features into a shared feature space. 2) Decomposed CZSL [20, 19, 40, 13, 14, 5, 8]: It disentangles visual features of simple primitives by designing two classifiers to predict state and object primitives respectively. Notably, recent breakthroughs in CZSL through CLIP-based methods (including both composed and decomposed methods) [30, 22, 23, 10]. These methods leverage frozen CLIP models to acquire textual embeddings of simple primitives, showcasing robust compositionality for zero-shot generalization. Nevertheless, they lack systematic investigation of the intrinsic relation between states and objects, ignoring the priority of visual cues. Specifically, Troika [10] adopts a Multi-Path framework to dynamically adjust prompts based on visual features, similar to the idea of the CoCoOP [42]. However, it may not fully leverage the inherent relationship between states and objects. In contrast, our PLO innovatively determines the observation order through language, allowing for a more nuanced understanding of state-object compositions and advancing the recognition of unseen combinations. This work represents the first effort, as far as we know, that models CZSL in a simple to complex manner by progressive language-based observations.
Large Language Models (LLMs) for CV Tasks. LLMs have ushered in a new era for natural language understanding and reasoning. There are two main types of solutions to integrate LLMs into CV: 1) Building multi-modal LLMs. Multi-modal pre-training aligns vision and language modalities in various ways, such as fine-tuning the visual encoder [35], training additional cross-attention layers [1], or reducing the size of extra layers and pre-trained lightweight modules [44, 17]. Such de facto paradigms require significant computing resources and training data, hindering further applications. 2) Deriving knowledge from LLMs. They query LLMs with meticulously curated prompts to obtain valuable information such as descriptions about category [18, 27, 31] or synthesized data [9]. By such means, vision tasks can enjoy powerful implicit knowledge within LLMs. In this paper, we harness the capabilities of GPT to generate observation prompts and facilitate compositional reasoning in CZSL. By leveraging the strengths of LLMs, PLO enhances the understanding of visual compositions and provides a novel approach to address CZSL challenges.
3 Progressive Language-based Observations
@@
Figure 2: (a) PLO-VLM: A two-step approach using a pre-observing classifier to dynamically determine the first observation. (b) PLO-LLM: A multi-step approach that observes composition-specific prompts from LLMs step-by-step.
@@
Formulation. Given an image $I$, CZSL focuses on the recognition of images belonging to composition categories $c\in C$, where the category set $C$ is defined as the Cartesian product of given state categories $S$ = $\{s_{1},\ldots,s_{|S|}\}$ and object categories $O$ = $\{o_{1},\ldots,o_{|O|}\}$, i.e., $C=S\times O$. During training, the models have access only to a set of seen compositions. In closed-world testing, the model must recognize images from both the seen compositions in $C(s)$ and the unseen compositions in $C(u)$, where $|C(s)\cup C(u)|\ll|C|$ during testing. In open-world testing, the model needs to recognize images from any composition within $C$. This scenario presents a more challenging task as the model must handle a broader range of compositions.
@@
In this section, we first introduce how to endow models with observing capabilities in Sec. 3.1. Then, we describe how to determine the step-by-step observation sequence for progressive comprehension, i.e., PLO-VLM and PLO-LLM, in Sec. 3.2 and Sec. 3.3, respectively.
3.1 CLIP-based Observing
@@
To enable observing ability, our PLO builds upon a pre-trained vision-language model: CLIP [34], which consists of an image encoder $En_{v}(\cdot)$ and a text encoder $En_{t}(\cdot)$ capable of mapping image features and semantic features from prompts (natural languages with category information, e.g., “a photo of ripe banana”) into a shared semantic space. By comparing the similarities between the two modalities in this shared space, we can discern whether language-based observations exist in a given image.
@@
1) Image Encoder $En_{v}(\cdot)$. We enhance the image encoder of CLIP with some lightweight parameter efficient fine-tuning (PEFT) strategies [6, 7], allowing it to effectively handle image features in CZSL. These PEFT strategies [^1] The detailed descriptions of the PEFT strategies are left in Appendix. allow the encoder to achieve performance levels comparable to full fine-tuning by transferring knowledge from CLIP while avoiding strong training biases with several learnable parameters.
@@
Specifically, the image encoder $En_{v}(\cdot)$ splits the input image into non-overlapping patches ($N$ in total) along with a pre-trained [CLS] token and positional embeddings, then generates a sequence of patch tokens. After that, the self-attention-based blocks, including the inserted learnable layers, are used to update the token sequence $\mathbf{{\mathcal{V}}}$ = $\{\mathbf{v}_{[\rm{CLS}]},\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{N}\}$. By optimizing the parameters of these inserted learnable layers during training while keeping the original image encoder frozen, PLO effectively incorporates primitive-specific information to improve observation capabilities. Finally, a linear layer is utilized to project the output [CLS] token $\hat{\mathbf{v}}_{[\rm{CLS}]}$, yielding the image representation $\mathbf{v}$ in the cross-modal shared space.
@@
2) Text Encoder $En_{t}(\cdot)$. Following [22], we employ soft prompt tuning, making prompts tokens better adapted to CZSL. Specifically, we formulate the prompt as a set comprising prefix context and category representations. By converting the prompt into learnable embeddings, we provide the model with the flexibility to adapt its language-based observation. The prompts of state, object, and composition, denoted as $\mathbf{P}_{s}$, $\mathbf{P}_{o}$, and $\mathbf{P}_{c}$, are formulated as:
$\mathbf{P}_{s}$ $=[\mathbf{x}_{0};\mathbf{x}_{1};\dots;\mathbf{x}_{M};\mathbf{x}_{s};\mathbf{o}],$ (1)
@@
where $[\mathbf{x}_{0};\dots;\mathbf{x}_{M}]$ represents the prefix context with the the word embedding of “a photo of” as an initialization, $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$ represent word embedding of the state and object categories, $\mathbf{o}$ represents word embedding of the word “object”. The context length is denoted by $M$. Similar to the token representation in the visual part, textual prompts are then converted to learnable embeddings and fed into the text encoder $En_{t}(\cdot)$ to update the patch tokens $\mathbf{{\mathcal{T}}}$ = $\{\mathbf{t}_{[\rm{CLS}]},\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{M}\}$ with self-attention blocks. Subsequently, the output $\hat{\mathbf{t}}_{[\rm{CLS}]}$ of the three types of prompts are fed into a single linear layer to project them as text representations $\mathbf{t}_{s}$, $\mathbf{t}_{o}$, and $\mathbf{t}_{c}$.
@@
Text representations $\mathbf{t}_{s}$, $\mathbf{t}_{o}$, and $\mathbf{t}_{c}$ and image representation $\mathbf{v}$ are then used in the subsequent steps of PLO-VLM and PLO-LLM process for compositional classification.
3.2 PLO-VLM

### v2

@@ Abstract.
Compositional zero-shot learning aims to recognize unseen state-object compositions by leveraging known primitives (state and object) during training. However, effectively modeling interactions between primitives and generalizing knowledge to novel compositions remains a perennial challenge. There are two crucial factors: large object-conditioned and state-conditioned variance, i.e., the appearance of states (or objects) can vary significantly when combined with different objects (or states). For instance, the state “old” can signify vintage design for a “car” or advanced age for a “cat”. In this paper, we argue that these variances can be mitigated by predicting composition categories based on salient observation cues. Therefore, we propose Progressive Language-based Observations (PLO), which can automatically determine the order of observation cues. These “observation cues” comprise a series of primitive concepts or graduated descriptions that allow the model to understand image content in a step-by-step manner. Specifically, PLO adopts pre-trained vision-language models (VLMs) to empower the model with observation capabilities. We further devise two variants: a two-step method (PLO-VLM) with a pre-observing classifier dynamically selecting the order of primitive concept-based cues, and a multi-step approach (PLO-LLM) using large language models (LLMs) to craft graduated description-based cues. Extensive tests on three datasets show PLO’s effectiveness in compositional recognition.
Compositional Zero-shot Learning, Large Language Models
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755208isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Object recognition
1. Introduction
@@
What enables us humans to recognize new concepts we have never encountered before? It comes down to our capacity to generalize learned knowledge to unseen domains (Mancini et al., 2021, 2022; Li et al., 2023a). For instance, when presented with concepts “green apple” and “yellow banana”, one can recognize and imagine the concept “green banana” by combining state “green” with object “banana”. Inspired by this innate cognitive ability of humans, Compositional Zero-Shot Learning (CZSL) emerges to tackle the challenge of recognizing unseen state-object compositions (e.g., “green banana”) by leveraging visible primitives (i.e., state and object) in compositional concepts during training and applying the learned knowledge during inference.
@@
Figure 1. Illustrations of challenges of CZSL and our PLO. (a) The challenge of object/state-conditioned variation: A perceptible variance emerges in the appearance of state/object primitives when juxtaposed in different compositions. (b) PLO-VLM/LLM: A two/multi-step observation approach dynamically controls the observation order for effective recognition.
@@
Effectively modeling the interactions between state and object primitives, as well as inducing the understanding of seen compositions to unseen ones, poses major challenges in CZSL (Mancini et al., 2021; Singh, 2022; Huo et al., 2024; Shao et al., 2025b). We argue that this task revolves around two key factors: 1) Object-conditioned Variance: Wherein the visual representations of the same state category can vary considerably when different objects are involved. As depicted in Figure 1(a), consider the state “old” in the context of modifying a “car” and a “cat”. For the “car”, it may refer to a vintage design characterized by classic curves and retro elements, conveying a sense of nostalgia and history. In contrast, for the “cat”, it denotes the advanced age of the feline, marked by features such as grey fur, which reflect the passage of time and the aging process. 2) State-conditioned Variance: It pertains to the variations in the appearance of an object when combined with different states. In Figure 1(a), for composition “peeled banana”, the “banana” exhibits a smooth texture and a pale appearance, as the outer peel is removed. In contrast, for “sliced banana”, the “banana” takes on a sliced appearance with visible segments. Previous CZSL methods often construct separate classifiers for recognizing states and objects simultaneously, overlooking their intrinsic relationship (Nayak et al., 2023; Lu et al., 2023a). Recent efforts have made strides in addressing the first factor by adopting a two-stage method with an object-then-state order (Wang et al., 2023; Huo et al., 2024; Kim et al., 2023). Prioritizing the object primitive prediction allows the model to capture salient visual cues (e.g., shapes), thereby enhancing the overall comprehension of compositions. Subsequently, armed with the knowledge of the object primitive, the CZSL model sequentially refines its understanding by classifying the state primitive conditioned on guided object features.
@@
Nonetheless, we argue that “all roads lead to Rome.”, and the human cognition process will progressively collect different observations for specific compositions, in a simple to complex manner (Konkle and Alvarez, 2024; Shi et al., 2024b; Bengio et al., 2009; Xiao et al., 2020; Li et al., 2023c). In certain cases, such as the composition “ripe banana” in Figure 1(b), the object, “banana”, possesses salient visual cues that make it easily recognizable due to its curving shape and vibrant yellow color. Once we establish that it is a “banana”, we can further analyze its state and recognize it as a “ripe banana” by observing additional visual cues, e.g., the presence of brown spots on the yellow skin. In contrast, compositions like “mashed banana” possess distinct visual features primarily related to the state “mashed” rather than the object. The mushy texture becomes the prominent aspect that captures our attention. Consequently, through analysis of extra visual features, e.g., yellow and sticky material, we can refine our recognition and discern it as a “mashed banana”.
@@
In this paper, inspired by the human-like cognition process, we propose a novel approach, Progressive Language-based Observations (PLO) for CZSL. Specifically, PLO automatically determines the order of progressive observation cues in the form of language, building upon the pre-trained vision-language models (VLMs), e.g., CLIP (Radford et al., 2021). These “observation cues” involve a series of primitive concepts or graduated descriptions that allow the model to observe the image’s content step by step. Due to training with image-text pairs, VLMs endow the model with observing capability by measuring the similarity between the two modalities within the same space. For automatic progressive observation, we propose two variants:
PLO-VLM is a two-step variant that adopts a pre-observing classifier based on VLM to dynamically determine the order of primitive concept-based cues based on image features. With the prior observed primitive knowledge (semantic features from primitive concepts, e.g., “banana”), we devise a cross-modal attention module that integrates this knowledge for category prediction of the remaining primitive. Since PLO-VLM requires fine-tuning with in-domain visual features, it can be well adapted to specific datasets.
PLO-LLM. While PLO-VLM advances CZSL, its observing classifier trained on specific datasets might not be generalizable enough for out-of-domain data (Zhu et al., 2023b; Wang et al., 2024) with distinct images. Thus, we further introduce PLO-LLM, evolving the concept into a more robust multi-step observation scheme. It employs large language models (LLMs) to directly generate domain-agnostic graduated descriptions (e.g., “yellow, mushy substance” in Figure 1(b)) for each composition category, thus obtaining a composition-specific observation order. This method allows us to extract salient features at each observation step according to universal descriptions, boosting the model’s ability to more robustly understand and recognize composition categories.
@@
Three prevalent CZSL datasets MIT-States (Isola et al., 2015), UT-Zappos (Naeem et al., 2021), and C-GQA (Yu and Grauman, 2014) are used for evaluation. Extensive results show that our PLO exceeds the current state-of-the-art CZSL methods with significant gains in both closed-world and open-world settings. In addition, cross-domain evaluation is conducted to demonstrate PLO-LLM’s excellent generalizability and robustness. In summary, the main contributions of our work are four-fold:
  • We argue that the concept of new compositions can vary significantly from the compositions in the dataset, which can be more easily recognized by the salient hints/cues from the concept of easily judged primitive or graduated descriptions from LLMs.
  • We propose an advanced mechanism for CZSL, automatically allocating the observation order in the form of primitive concepts or graduated descriptions, enabling effective prediction of unseen state-object compositions.
@@
  • We introduce two variants: 1) PLO-VLM: using a pre-observing classifier to dynamically determine observation cue order based on image features. 2) PLO-LLM: employing LLMs to directly devise composition-specific graduated descriptions for observation.
@@
  • Extensive results on multiple datasets demonstrate that our PLO outperforms existing CZSL methods, showcasing its effectiveness in recognizing unseen state-object compositions.
2. Related Work
@@
Compositional Zero-Shot Learning (CZSL). CZSL is a prominent research area, and state-of-the-art methods can be broadly categorized into two directions: 1) Composed CZSL (Misra et al., 2017; Purushwalkam et al., 2019; Naeem et al., 2021; Anwaar et al., 2022; Mancini et al., 2022): It focuses on directly classifying compositional state-object, which projects whole compositional visual features into a shared feature space. 2) Decomposed CZSL (Li et al., 2020, 2022b; Zhang et al., 2022a; Karthik et al., 2022; Khan et al., 2023; Hao et al., 2023; Hu and Wang, 2023; Jiang and Zhang, 2024): It disentangles visual features of simple primitives by designing two classifiers to predict state and object primitives respectively. Notably, recent breakthroughs in CZSL through CLIP-based methods (including both composed and decomposed methods) (Nayak et al., 2023; Lu et al., 2023a, b; Huang et al., 2023; Zheng et al., 2024). These methods leverage frozen CLIP models to acquire textual embeddings of simple primitives, showcasing robust compositionality for zero-shot generalization. Nevertheless, they lack systematic investigation of the intrinsic relation between states and objects, ignoring the priority of visual cues. Specifically, Troika (Huang et al., 2023) adopts a Multi-Path framework to dynamically adjust prompts based on visual features, similar to the idea of the CoCoOP (Zhou et al., 2022a). However, it may not fully leverage the inherent relationship between states and objects. In contrast, our PLO innovatively determines the observation cue order through language, allowing for a more nuanced understanding of state-object compositions and advancing the recognition of unseen combinations. This work represents the first effort, as far as we know, that models CZSL in a simple-to-complex manner by progressive language-based observations.
@@
Figure 2. (a) PLO-VLM: A two-step approach using a pre-observing classifier to dynamically determine the first observation. (b) PLO-LLM: A multi-step approach that observes composition-specific graduated descriptions from LLMs step-by-step.
Large Language Models (LLMs) for CV Tasks. LLMs have ushered in a new era for natural language understanding and reasoning. There are two main types of solutions to integrate LLMs into CV: 1) Building multi-modal LLMs. Multi-modal pre-training aligns vision and language modalities in various ways (Zang et al., 2024; Shi et al., 2024a; Liu et al., 2024; Zhu et al., 2024; Ming and Li, 2024; Wang et al., 2025), such as fine-tuning the visual encoder (Tsimpoukelli et al., 2021), training additional cross-attention layers (Alayrac et al., 2022), or reducing the size of extra layers and pre-trained lightweight modules (Zhu et al., 2023a; Li et al., 2022a). Such de facto paradigms require significant computing resources and training data, hindering further applications. 2) Deriving knowledge from LLMs. They query LLMs with meticulously curated prompts to obtain valuable information such as descriptions about category (Li et al., 2023b; Menon and Vondrick, 2023; Novack et al., 2023; Zhang et al., 2023; Yang et al., 2023) or synthesized data (Hu et al., 2022a). By such means, vision tasks can enjoy powerful implicit knowledge within LLMs. In this paper, we harness the capabilities of GPT to generate graduated descriptions and facilitate compositional reasoning in CZSL. By leveraging the strengths of LLMs, PLO enhances the understanding of visual compositions and provides a novel approach to address CZSL challenges.
3. Approach
@@
Formulation. Given an image $I$, CZSL focuses on the recognition of images belonging to composition categories $c\in C$, where the category set $C$ is defined as the Cartesian product of given state categories $S$ = $\{s_{1},\ldots,s_{|S|}\}$ and object categories $O$ = $\{o_{1},\ldots,o_{|O|}\}$, i.e., $C=S\times O$. During training, the models have access only to a set of seen compositions. In closed-world testing, the model must recognize images from both the seen compositions in $C(s)$ and the unseen compositions in $C(u)$, where $|C(s)\cup C(u)|\ll|C|$. In open-world testing, the model needs to recognize images from any composition within $C$. This scenario presents a more challenging task as the model must handle a broader range of compositions.
@@
In this section, we first introduce how to endow models with observing capabilities in Sec. 3.1, then describe how to determine the order of step-by-step observation cues for progressive comprehension, i.e., PLO-VLM and PLO-LLM, in Sec. 3.2 and Sec. 3.3.
3.1. CLIP-based Observing
@@
To enable observing ability, PLO builds upon a pre-trained vision-language model: CLIP (Radford et al., 2021), which consists of an image encoder $En_{v}(\cdot)$ and a text encoder $En_{t}(\cdot)$ capable of mapping image features and semantic features from prompts (natural languages with category information, e.g., “a photo of ripe banana”) into a shared semantic space. By comparing the similarities between the two modalities in this shared space, we can discern whether language-based observations exist in a given image.
@@
1) Image Encoder $En_{v}(\cdot)$. We enhance the image encoder of CLIP with some lightweight parameter efficient fine-tuning (PEFT) strategies (Houlsby et al., 2019; Hu et al., 2022b; Zheng et al., 2024) to effectively handle image features in CZSL. These PEFT strategies allow the encoder to achieve performance levels comparable to full fine-tuning by transferring knowledge from CLIP while avoiding strong training biases with several learnable parameters. Details of the PEFT strategies are in the Appendix.
@@
Specifically, $En_{v}(\cdot)$ splits the input image into non-overlapping patches ($N$ in total) along with a pre-trained [CLS] token and positional embeddings, then generates a sequence of patch tokens. After that, the self-attention-based blocks, including the inserted learnable layers, are used to update the token sequence $\mathbf{{\mathcal{V}}}$ = $\{\mathbf{v}_{[\rm{CLS}]},\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{N}\}$. By optimizing the parameters of inserted learnable layers during training while keeping the original image encoder frozen, PLO effectively incorporates primitive-specific information to improve observation capabilities. Finally, a linear layer is used to project the output [CLS] token $\hat{\mathbf{v}}_{[\rm{CLS}]}$, yielding the image representation $\mathbf{v}$ in the cross-modal shared space.
@@
2) Text Encoder $En_{t}(\cdot)$. Following (Lu et al., 2023a), we employ soft prompt tuning, making prompt tokens better adapted to CZSL. Specifically, we formulate the prompt as a set comprising prefix context and category representations. By converting the prompt into learnable embeddings, we provide the model with the flexibility to adapt its language-based observation. The prompts of state, object, and composition, denoted as $\mathbf{P}_{s}$, $\mathbf{P}_{o}$, and $\mathbf{P}_{c}$, are formulated as:
(1) $\mathbf{P}_{s}$ $=[\mathbf{x}_{0};\mathbf{x}_{1};\dots;\mathbf{x}_{M};\mathbf{x}_{s};\mathbf{o}],$
@@
where $\mathbf{x}_{0};\dots;\mathbf{x}_{M}]$ denotes the prefix context with the word embedding of “a photo of” as initialization, $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$ represent word embedding of the state and object categories, $\mathbf{o}$ is word embedding of the word “object”. The context length is denoted by $M$. Similar to the token representation in the visual part, these textual prompt embeddings (i.e., context embeddings, category embeddings, and word embedding $\mathbf{o}$) are then converted to learnable embeddings and fed into the text encoder $En_{t}(\cdot)$ to update the patch tokens $\mathbf{{\mathcal{T}}}$ = $\{\mathbf{t}_{[\rm{CLS}]},\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{M}\}$ with self-attention blocks. These learnable embeddings are shared across the three prompts. Subsequently, the output $\hat{\mathbf{t}}_{[\rm{CLS}]}$ of the three types of prompts are fed into a single linear layer to project as text representations $\mathbf{t}_{s}$, $\mathbf{t}_{o}$, and $\mathbf{t}_{c}$.
@@
These text and image representations are then used in subsequent PLO-VLM and PLO-LLM steps for compositional classification.
3.2. PLO-VLM